OpenStudy (anonymous):

It take 9hrs for 6 cooks to prepare the food for a wedding rehearsal dinner. How long will it take for 2 cooks to prepare the dinner. What kind of variation applies to this situation Is this Inverse or direct.

OpenStudy (amistre64):

it its an inverse relation; the more cooks the less time it takes

OpenStudy (amistre64):

inverses go in opposite directions; more = less and less = more

OpenStudy (amistre64):

9 = k/6 k = 9.6 = 54 h = 54/2 = 27